OpenStudy (anonymous):

How to evaluate the integral of:

OpenStudy (anonymous):

\[\int\limits_{0}^{1} \sin(3\pi(t))dt\]

OpenStudy (anonymous):

I know hot to use FTC and U substitution, but I'm having trouble with knowing when to use which..

OpenStudy (anonymous):

anti derivative is easy enough

OpenStudy (anonymous):

*how

OpenStudy (anonymous):

\[-\frac{1}{3\pi}\cos(3\pi t)\]

OpenStudy (anonymous):

a mental u-sub

OpenStudy (anonymous):

Where did the (1/3pi) come from? I get why it's negative cos(...)

OpenStudy (anonymous):

what is the derivative of \[-\cos(3\pi t)\]

OpenStudy (anonymous):

sin(3(pi)t)?

OpenStudy (anonymous):

oh no !

OpenStudy (anonymous):

you need the chain rule for that one

OpenStudy (anonymous):

OH does this involve chain rule?

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

let me know when you get \[3\pi \sin(3\pi t)\]

OpenStudy (anonymous):

yeah, I got that!

OpenStudy (anonymous):

but that isn't what you want is it? you just want \[\sin(3\pi t)\]

OpenStudy (anonymous):

Oh okay, I understand now. So in this case, would u just be the entire function? sin(3pi(t))?

OpenStudy (anonymous):

u shmoo you don't want \[3\pi \sin(3\pi t)\] you just want \[\sin(3\pi t)\] so divide ! make it \[-\frac{1}{3\pi}\cos(3\pi t)\] and it will work out nicely it is a u - sub you do in your head

OpenStudy (anonymous):

Okay. So generally, how can I look at an integral and know which method to use?